A store manager wishes to reduce the price on his fresh ground coffee by mixing 2 grades. If he has 50 pounds of coffee which sells for $5 per pound, how much coffee worth $3 per pound must he mix with it so that he can sell the final mixture for $4.25 per pound?
Answer by checkley77(12569)

50*5+3x=4.25(50+x)
250+3x=212.5+4.25x
3x-4.25x=212.5-250
-1.25x=-37.5
x=-37.5/-1.25
x=30 pounds of $3 coffee is required.
Proof:
50*5+3*30=4.25(50+30)
250+90=4.25*80
340=340